OpenStudy (anonymous):

2+4+6+...+100. Use Gauss's approach to find the sum.

OpenStudy (anonymous):

multiply the number of terms, 50, by the average of the first and last terms, 2 and 100.\[n(\frac{ a+b }{ 2 })\]n=50, a=2, and b=100. So the sum is 50 times 102/2, or 50*51=2550

OpenStudy (anonymous):

thanks